OpenStudy (anonymous):

solve the following simultaneous equations. (i) 3x+y=6 y=x-2 ii)3x=5y=-6 x+2y=3

OpenStudy (ikram002p):

3x+y=6 y=x-2 3x+(x-2)=6 4x=8 x= :P

OpenStudy (ikram002p):

3x=5y=-6 x+2y=3 multiply both the second equation by 3 :P 3x+5y=-6 3x+6y=9 -------- 0x-y=-12 y=12

OpenStudy (anonymous):

y=0 x=2

OpenStudy (anonymous):

y=15 x=-27

OpenStudy (anonymous):

thanks guys both of your answer are correct. i wish i could give two medal :( ....maybe one medal for ikram and fan for sargams :) thanks guys really appreciate the help.....